OpenStudy (anonymous):

WILL MEDAL AND FAN PLEASE HELP LOVELIES xxxx Which numbers are greater than 7 ������ 10���������4? Choose all answers that are correct. A. 8 x10���4 B. 9 ������ 10���������5 C. 4 ������ 10���������4 D. 3 ������ 10���������2 E. 7 ������ 10���������5

OpenStudy (anonymous):

A and B. Solve it!

OpenStudy (anonymous):

I thought it was b and c??

OpenStudy (anonymous):

thankyou btw :)

OpenStudy (anonymous):

Np, thanks. C isn't because if you solve it, the answer is 36...

OpenStudy (solomonzelman):

wowow

OpenStudy (solomonzelman):

Let \(\Large\color{red}{ N_1=a\times 10^c }\) \(\Large\color{slateblue}{ N_2=x\times 10^z }\) ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ where "a" and "x" are integer values that satisfy the statement \(\Large\color{slateblue}{ 0>|s|>10 }\) "c" and "z" are just some integers. this is just definition I will use in my explanation

OpenStudy (solomonzelman):

oh, sorry "a" and "x" can be any rational numbers

OpenStudy (solomonzelman):

\(\Large\color{red}{ N_1=a\times 10^c }\) \(\Large\color{slateblue}{ N_2=x\times 10^z }\) and, in order for \(\Large\color{red}{ N_1\color{black}{>} }\) \(\Large\color{slateblue}{ N_2 }\) 1 of the following must be true 1) in a case where \(\Large\color{red}{ c\color{black}{=} }\)\(\Large\color{slateblue}{ z}\), must be so that \(\Large\color{red}{ a\color{black}{>} }\)\(\Large\color{slateblue}{ x}\). 2) (or, just) \(\Large\color{red}{ c\color{black}{>} }\)\(\Large\color{slateblue}{ z}\).